OpenStudy (anonymous):

Simplify the expression (x^4)^3/5

OpenStudy (johnweldon1993):

since the expression is raised to a power...you would multiply these exponents.. x^4 times ^3/5 what is 4 times 3/5?

OpenStudy (anonymous):

\[\frac{ 1 }{4^{-4} }\]

OpenStudy (anonymous):

12/5?

OpenStudy (johnweldon1993):

12/5 is correct so you have x^12/5

OpenStudy (anonymous):

What about the one I 'drew'?

OpenStudy (johnweldon1993):

oh...didn't even see that lol well okay...this will be a complex fraction so how do we write 4^-4? ...you would write 1/4^4 right? so we'll all together have \[\frac{ 1 }{ \frac{1}{ 4^4 } }\] how do we solve this? we "FLIP" the bottom fraction...and change the whole thing to multiplication so it would be 1 over 4^4 over 1 which would become 1 times 4^4 over 1 so what is 1 times 4^4?

OpenStudy (anonymous):

256

OpenStudy (johnweldon1993):

right....and 256/1....yeah 256 lol...that's what your answer would be

OpenStudy (anonymous):

Oh, okay. Thanks a lot, man.

OpenStudy (johnweldon1993):

no prob bud!